### v1

# Why iCloud Fails:

Abstract
iCloud Drive presents a filesystem interface but implements cloud synchronization semantics that diverge from POSIX in fundamental ways. This divergence is not an implementation bug; it is a Category Mistake—the same one that pervades distributed computing wherever Forward-In-Time-Only (FITO) assumptions are embedded into protocol design. Parker et al. showed in 1983 that network partitioning destroys mutual consistency; iCloud adds a user interface that conceals this impossibility behind a facade of seamlessness. This document presents a unified analysis of why iCloud fails when composed with Time Machine, git, automated toolchains, and general-purpose developer workflows, supported by direct evidence including documented corruption events and a case study involving 366 GB of divergent state accumulated through normal use. We show that the failures arise from five interlocking incompatibilities rooted in a single structural error: the projection of a distributed causal graph onto a linear temporal chain. We then show how the same Category Mistake, when it occurs in network fabrics as link flapping, destroys topology knowledge through epistemic collapse. Finally, we argue that Open Atomic Ethernet (OAE) transactional semantics—bilateral, reversible, and conservation-preserving—provide the structural foundation for resolving these failures, not by defeating physics, but by aligning protocol behavior with physical reality.
1  Introduction: iCloud Is Not a Filesystem
Users are encouraged to believe that iCloud Drive is a transparent extension of the local filesystem. Apple’s marketing promises seamless synchronization: files appear on every device, changes propagate automatically, and the experience is indistinguishable from local storage.
This promise routinely fails to deliver. Not because Apple’s engineers are incompetent, but because the promise itself rests on a Category Mistake.
A filesystem provides:
  • stable paths that do not change based on operational state,
  • atomic operations (rename, create, link) with POSIX semantics,
  • advisory locking that coordinates concurrent access,
  • consistent directory listings that reflect recent writes,
  • the guarantee that listed files are present and readable.
iCloud Drive can violate each of these properties under documented conditions:
  • File paths change based on sync state (legacy .icloud stubs; modern dataless files),
  • Operations are not atomic across the sync boundary,
  • NSFileCoordinator replaces POSIX locking with optional, cooperative IPC,
  • Directory listings include dataless files with no data extents,
  • Files may be listed but require network fetches before content is available.
iCloud Drive is a distributed negotiation system that presents a filesystem facade. Treating this facade as authoritative is the Category Mistake that produces the failures users experience daily.
The consequences extend beyond inconvenience. As Dominik Mayer documented, iCloud Drive silently deletes user content when version conflicts arise [23]. Mayer identifies the core of the problem as a product decision: “the core issue is the product decision to not be transparent about version conflicts. There is no central interface to manage them, nor are they somehow represented on the hard drive so you could deal with them whenever you are looking at your files.” The system does not merely fail to synchronize correctly; it actively conceals the fact that it has failed.
This is not a new problem. Parker et al. established in 1983 that network partitioning can “completely destroy mutual consistency” in replicated filesystems [24]. The theoretical impossibility was clear forty years ago. What iCloud adds is a user interface that hides the impossibility behind a facade of seamlessness, converting a known hard problem into silent data destruction.
2  The Structural Thesis
The failures of iCloud, when composed with Time Machine, git, automated toolchains, or any program that assumes POSIX semantics, are not random bugs. They are projection artifacts caused by a single structural error:
Projecting a distributed causal graph onto a linear temporal chain destroys essential structure. Information is lost. The losses manifest as corruption, conflicts, stalls, and silent data destruction.
This is an instance of Forward-In-Time-Only (FITO) thinking [1]. FITO replaces partial order with total order and treats the result as authoritative history. Cloud sync systems operate in causal order. Snapshot systems, timestamp-based conflict resolution, and lock-file protocols operate in wall-clock order. These are not equivalent.
Parker et al. showed in 1983 that mutual consistency cannot be maintained under network partitioning [24]. Every cloud sync system—iCloud, Google Drive, Dropbox, OneDrive—operates in exactly this regime: intermittently connected devices that partition whenever a laptop lid closes or a phone enters a tunnel. The failures documented below are not edge cases. They are the predicted consequences of a problem the distributed systems community identified and characterized four decades ago.
3  The Architecture of iCloud Drive
To understand why iCloud fails, we must understand what it actually is.
3.1  The Coordination Layer
iCloud Drive does not use POSIX advisory locks. It employs NSFileCoordinator, a higher-level coordination mechanism that mediates access between applications and the bird sync daemon via inter-process communication [9]. This coordination is cooperative and optional: any process can bypass it by operating directly on files. The sync daemon itself has exclusive access during upload/download operations, producing the notorious “Operation not permitted” errors when tools like rm attempt to modify syncing files.
3.2  Temporal Semantics: Last-Writer-Wins
iCloud resolves conflicts using modification timestamps—last-writer-wins. It does not use vector clocks or logical timestamps [10]. This means:
  • clock skew across devices produces false ordering,
  • “later” is treated as “more correct,”
  • a single linear history is assumed where none exists.
None of these assumptions hold reliably in a distributed, intermittently connected environment.
3.3  CloudKit and FoundationDB
The server side runs on CloudKit, which is built on FoundationDB’s Record Layer—an ACID-compliant, multi-tenant distributed database [12]. The server thus appears to provide strong transactional semantics. The failures occur at the client, where the sync daemon projects this rich distributed state onto a local filesystem that cannot represent it. The irony is structural: ACID guarantees on the server are erased by the projection onto the client.
3.4  The fileproviderd Daemon
Since macOS High Sierra, the actual work of iCloud Drive synchronization is performed by fileproviderd, a system daemon that manages the File Provider framework [25]. This daemon handles metadata synchronization, on-demand file access through placeholders, and conflict resolution for both iCloud and third-party providers (Dropbox, OneDrive).
The daemon exhibits several documented failure modes under load:
  • Files hang indefinitely in “Uploading” state during large sync operations,
  • CPU and memory consumption spikes during bulk synchronization,
  • Error propagation from CloudKit is incomplete—applications receive no notification when server-side operations fail,
  • Rapid multi-device edits produce merge conflicts that the daemon resolves by creating duplicates or silently discarding one version.
The daemon’s behavior can be observed via log stream --predicate ’subsystem == "com.apple.FileProvider"’, but the logs reveal an eventual-consistency model with no mechanism for applications to verify that their writes have been durably committed to the cloud. The daemon operates as a best-effort system that presents itself as a reliable one.
3.5  Dataless Files and the Eviction Problem
Since macOS Sonoma, iCloud Drive uses APFS dataless files: file metadata and extended attributes are stored locally, but data extents are absent [13]. A dataless file:
  • is listed by ls and returns valid metadata from stat(),
  • has no data blocks allocated,
  • triggers an implicit network fetch when read,
  • is indistinguishable from a real file without inspecting the SF_DATALESS flag.
This breaks the expectation—implicit in decades of POSIX practice, if not in the formal specification—that listed files are present and readable. Tools that rely on this expectation—including git, make, xcodebuild, and AI coding agents—fail unpredictably.
4  Incompatibility 1: Time Machine
Time Machine archives filesystem trees indexed by time. iCloud Drive negotiates distributed event graphs. A tree snapshot is not equivalent to a distributed agreement state.
4.1  The Projection
Time Machine defines an implicit projection:
$F:\text{Sync Category}\rightarrow\text{Snapshot Category}$
It forgets the event graph and retains only the instantaneous local tree. This projection is not faithful: distinct distributed histories can map to the same local snapshot; the same distributed history can map to different snapshots depending on timing [2].
4.2  What Is Lost
  • Concurrency lineage
  • Device-of-origin
  • Conflict sets
  • Merge provenance
  • Hydration status (dataless vs. materialized)
4.3  The Sonoma Cascade
macOS Sonoma demonstrated this incompatibility dramatically: multiple users reported Time Machine backups failing to complete while iCloud Drive was actively syncing. The reported root cause was a circular dependency: Time Machine requires a stable filesystem state to snapshot; iCloud’s sync daemon continuously modifies the filesystem; the FSEvents database does not settle; and Time Machine cannot acquire a coherent change set [15].
4.4  Restore Semantics
When a Time Machine backup is restored while iCloud is active, the sync daemon may treat cloud state as authoritative and overwrite the restored local state [11]. The “correct past” that Time Machine preserved can be destroyed by the subsequent sync. This is not a bug in either system. It is a structural consequence of two systems with incompatible models of truth.
5  Incompatibility 2: Git
Git depends on POSIX filesystem semantics at a level of detail that iCloud cannot provide.
5.1  Git’s Locking Model
Git implements mutual exclusion through lock files created with open(O_CREAT|O_EXCL)—an atomic create-if-not-exists operation. The write-sync-rename pattern (write $\to$ fsync $\to$ rename) provides transactional semantics: either the update completes fully, or the repository is unchanged [16].
5.2  How iCloud Defeats Git
  1. 1.
    Lock file propagation: iCloud treats .git/index.lock as an ordinary file and syncs it to other devices. A device receiving this file believes another process holds the lock and refuses to operate.
  2. 2.
    Ref corruption via numbered suffixes: When iCloud detects concurrent modifications to refs/heads/main, it renames one version to refs/heads/main 2. Git never looks for numbered suffixes. The commit is orphaned and invisible.
  3. 3.
    Packfile tearing: Because iCloud syncs files independently, a .pack file and its corresponding .idx file can arrive at a remote device in inconsistent states. If the packfile is still being written when sync begins, the result is an index whose offsets no longer match the pack contents.
  4. 4.
    Race conditions: Git’s rename() of index.lock to index races with iCloud’s sync of the lock file itself. The sync daemon can observe and propagate intermediate states that git intends to be invisible.
5.3  Observed Failures
These failure modes are not hypothetical. Multiple instances of iCloud-induced git corruption have been directly observed during the preparation of this document and related work, including: iCloud propagating intermediate file states during LaTeX collaboration; silent filename swapping that undid deliberate file creation; and the sync daemon seizing .git/index.lock within seconds of git init, rendering a repository permanently inoperable with “Operation not permitted” on every file. These incidents are documented with verbatim error output in Appendix A.
5.4  The Bundle Workaround
The established workaround is to store git repositories outside iCloud and export .bundle files for archival [8]. A bundle is a single file; iCloud’s conflict resolution (numbered suffixes) produces two complete bundles rather than a corrupted repository tree. The workaround succeeds precisely because it removes git from the cloud sync system entirely. This is an admission that the two systems are structurally incompatible.
6  Incompatibility 3: Automated Toolchains
Any automated toolchain—build systems, continuous integration pipelines, scripting frameworks, coding agents—assumes it is operating on a local POSIX filesystem with strong consistency, immediate write visibility, and no external modification of files between operations. The following failure modes are logical consequences of composing such a toolchain with iCloud’s sync semantics. Some have been directly observed; others follow from the architectural mismatch documented in the preceding sections.
6.1  The Toolchain-Sync Race
When an automated toolchain edits a file in an iCloud-synced directory, the following race is structurally possible:
  1. 1.
    The agent writes an incomplete intermediate state.
  2. 2.
    The sync daemon detects the change and begins uploading.
  3. 3.
    The agent completes the edit.
  4. 4.
    The sync daemon may have already propagated the incomplete state.
The consequence: other devices can receive corrupted files. The agent itself may read stale or dataless files on subsequent operations.
6.2  File Watcher Failure
Claude Code’s atomic write operations (write to temporary file, then rename) have been observed to fail to trigger filesystem event notifications that tools such as watchman expect [19]. Independently, iCloud Drive’s own interaction with file-watcher libraries is known to produce unreliable event streams [20]. The combination means that neither the agent nor the sync daemon can maintain reliable knowledge of the filesystem’s current state.
6.3  The Placeholder Trap
When “Optimize Mac Storage” is enabled, files that an agent successfully read in one session may be evicted by iCloud between sessions. On the next invocation, the agent encounters dataless files: stat() succeeds, but read() triggers a network fetch that may time out, return stale data, or fail entirely. The filesystem has changed state underneath the agent without any operation by the agent. This is a direct consequence of the dataless file architecture described in Section 3.
7  Incompatibility 4: The Filesystem Dialect Problem
Beyond the POSIX violations, iCloud introduces semantic loss at the metadata layer:
  • Extended attributes marked with Apple’s #S suffix are preserved during sync; others are silently stripped.
  • The com.apple.ResourceFork extended attribute is typically stripped during sync, even with the #S suffix [14].
  • Case sensitivity, path normalization, and forbidden characters differ across platforms (macOS, iOS, Windows via iCloud for Windows), creating mismatches invisible at the byte level.
  • APFS clones—which share data blocks on the originating device—appear as full copies on other devices. The clone semantics do not replicate.
These losses are individually minor but cumulatively devastating. A project that depends on extended attributes, resource forks, or case-sensitive paths will experience silent corruption that no error message reveals.
8  Case Study: The 366 GB Archive
The preceding incompatibilities are not isolated incidents. Over years of use, they compound into an operational crisis. One of the authors accumulated, through normal use of iCloud Drive across multiple devices, a 366 GB “iCloud Drive (Archive)” folder containing 110 top-level items that had diverged from the 405 GB primary iCloud Drive.
Analysis revealed:
  • 89 folders existed only in the archive—files that iCloud had silently removed from the primary store.
  • 57 folders existed in both locations with divergent contents.
  • Within the shared folders, individual files with identical names had different MD5 checksums—silent version drift with no conflict notification.
  • The FIGURES directory alone contained 1,406 files unique to the archive that had vanished from the primary store.
Resolving this required writing custom Python scripts to perform MD5-based comparison of every file in both trees, categorize conflicts, and merge without data loss [28]. The scripts could not rely on timestamps (iCloud’s conflict resolution had already corrupted them) and instead had to treat file content as the only source of truth.
The same pattern appeared elsewhere in the filesystem: a personal document collection had been duplicated into multiple parallel folder hierarchies by iCloud sync, requiring MD5-based de-duplication to determine which copies were authoritative [29].
This is what the Category Mistake looks like at operational scale. A system that promises seamless synchronization instead produced hundreds of gigabytes of unresolvable divergence, silent deletions, and phantom duplicates—requiring weeks of manual effort and custom tooling to repair.
9  The Deeper Pattern: FITO in Sync Protocols
The preceding four incompatibilities share a common root: all arise from Forward-In-Time-Only assumptions embedded in iCloud’s sync protocol.
9.1  Timestamp Primacy
Conflicts are resolved by modification time. This assumes clocks are meaningful across devices, that “later” implies “more correct,” and that a single linear history exists. None of these hold.
9.2  Intermediate State Leakage
Application save protocols rely on multi-step sequences intended to be atomic. iCloud’s sync engine propagates intermediate states, exposing partially written files. This is a direct consequence of replaying a forward timeline rather than reconciling intended state.
9.3  Smash-and-Restart Recovery
When client-side state diverges, iCloud resorts to reset-and-rescan: discard causal structure, crawl forward again. This approach is fragile, non-idempotent, and produces the duplicate trees, resurrected deletions, and silent overwrites that users report.
9.4  Non-Idempotent Retry Semantics
Retries duplicate intent rather than reconcile it, producing multiple versions of what was meant to be a single logical object. This is timeout-and-retry (TAR) at the filesystem layer—the same pattern that causes retry storms in network protocols [1].
10  Empirical Evidence: Network Partitions Cause Application Failures
The structural pattern described above—FITO assumptions converting transient uncertainty into catastrophic, silent, and permanent failures—is not confined to file synchronization. It has been empirically confirmed across a wide range of distributed systems by the University of Waterloo group.
Alquraan et al. conducted a comprehensive study of 136 system failures attributed to network-partitioning faults across 25 widely used distributed systems [33]. Their findings are striking:
  • 80% of the failures had catastrophic impact on the system.
  • Data loss was the most common consequence (26.6% of failures), followed by reappearance of deleted data, broken locks, and system crashes.
  • 90% of the failures were silent—the system did not raise an alert, write a log entry, or return an error to the client.
  • 21% led to permanent damage that persisted even after the network partition healed.
  • The majority of failures required little or no client input, could be triggered by isolating a single node, and were deterministic.
The extended study by Alkhatib et al. on partial network partitions—where some but not all nodes lose connectivity—found the same pattern in 13 popular systems [34]. The failures were catastrophic, easily manifested, and mainly due to design flaws in core system mechanisms including scheduling, membership management, and ZooKeeper-based configuration management.
These results are directly relevant to iCloud Drive. iCloud operates in exactly the regime these studies characterize: intermittent connectivity between devices (a laptop lid closes, a phone enters a tunnel, an airplane enters flight mode) creates network partitions that the sync protocol must handle. The Waterloo findings confirm that the consequences are not merely theoretical. Silent data loss, permanent damage, and undetected divergence are the empirically measured outcomes of partition-handling failures across the industry’s most widely deployed systems.
10.1  The Limits of Overlay Solutions
The Waterloo group proposed Nifty, a transparent overlay network layer that masks partial partitions by rerouting packets around failed links [34]. While the empirical diagnosis is precise, the proposed solution illustrates a characteristic limitation: it operates above the link layer (L3/L4 overlay), attempting to compensate for failures that originate at or below L2.
This is the same architectural pattern that produces iCloud’s failures: applying higher-layer workarounds to lower-layer problems. Overlay rerouting cannot address the fundamental issue—that the link itself has entered an ambiguous state where delivery is unknown—because the overlay has no access to the bilateral link state. The overlay can detect that packets are not arriving and reroute them, but it cannot determine whether the original packets were delivered, partially delivered, or lost. It therefore introduces the same three-state ambiguity (delivered, not delivered, unknown) that FITO protocols collapse into error.
As demonstrated by the OAE link protocol (Section 13), resolving partition-induced ambiguity requires bilateral state at the link boundary itself. The fix must be at L2, not above it, because that is where the causal information about delivery exists. No amount of overlay routing can recover information that was never captured at the point of uncertainty.
10.2  The Broken Link and the Triangle Topology
There is a deeper objection that must be addressed. If a link is truly broken, no information can flow across it in either direction. Bilateral state at the endpoints is necessary but not sufficient: both sides know the link is down (disconnection is the one failure hazard that a direct link can unambiguously detect [30]), but neither side can communicate this knowledge to the other.
This is where the distinction between links and switched networks becomes critical. In a switched network, the failure modes are numerous and opaque: packets are silently dropped, delayed, duplicated, and reordered by intermediate switches. The endpoints cannot distinguish a slow node from a dead one, a congested path from a partitioned one. The ambiguity is irreducible because the intermediate hops destroy causal information.
A direct link has exactly one failure hazard: disconnection. And disconnection is straightforward to detect—the absence of the expected token is unambiguous. Both endpoints know the link is down. What they lack is a path to each other through which to coordinate recovery.
The resolution requires a triangle topology: a minimum of three nodes connected by direct links, such that when any single link fails, the remaining two links provide an alternative path. This is the minimal causal structure in which shared state can be maintained despite a single-link failure. One round trip (Alice $\to$ Bob $\to$ Alice) establishes the triangle inequality—the minimal geometric structure that supports bilateral knowledge [31].
This is precisely what overlay networks attempt to provide, but at the wrong layer. An overlay reroutes packets around a failed switched path, but the rerouted path still traverses switches that can silently corrupt the causal chain. A triangle of direct OAE links provides the same topological redundancy with the causal guarantees intact: each link in the triangle maintains bilateral state, and the failure of any one link is communicated via the other two with no loss of causal information.
Figure 1: A chain of nodes connected by AELinks. Charlie (below) forms a triangle with Alice and Bob, providing a recovery path when the Alice–Bob link fails. From [32].
Figure 2: Triangle failover as a three-party transaction. Charlie acts as the transaction manager (TM) for the Alice–Bob link, coordinating a 2PC recovery via the surviving links. Every triangle is a recovery resource; the role of TM rotates symmetrically among all three nodes. From [32].
In OAE terminology, each link forms part of a three-node consensus group—a Reliable Link Failure Detector (RLFD) that operates silently at L2 [32]. Every triangle is a recovery resource. Consider a chain of nodes: Alice–Bob–Carol, connected by direct AELinks (Figure 1). A third node, Charlie, sits off to the side of the chain, with direct links to both Alice and Bob. When the Alice–Bob link fails, both endpoints detect the disconnection unambiguously—and Charlie, as the transaction manager (TM), can coordinate recovery via its surviving links to each of them. This is local failover without timeout-and-retry: the triangle structure provides the alternative path and the causal guarantees simultaneously.
The combinatorial argument makes the advantage concrete. In a switched network of just four nodes, each pair connected through a Clos fabric, Colm MacCárthaigh has observed that with n(n1)/2=6n(n-1)/2=6 logical links and four possible states per link (both directions up, each direction down independently, both down), the system has $4^{6}-1=4{,}095$ possible failure configurations [32]. Link failures in a Clos are invisible to the endpoints—the switch fabric silently drops, delays, duplicates, or reorders packets without notification. In a direct-link mesh, every one of those failure modes is immediately visible to the affected endpoints. Triangle supervision eliminates silent failures.
The triangle failover mechanism is, in fact, a three-party transaction (Figure 2). When the link between Alice and Bob fails, Charlie—positioned off the chain—acts as the transaction manager (TM) and initiates recovery on the Alice–Bob link by harvesting events from its own ports to Alice and Bob. This is exactly the two-phase commit protocol (2PC) that Jim Gray described: a coordinator proposes, the participants vote, and the coordinator resolves [35, 36]. The symmetry of the triangle makes the role assignment natural and complete: Alice is the TM for link Bob–Charlie, Bob is the TM for link Charlie–Alice, and Charlie is the TM for link Alice–Bob. Any node that detects a failure on one of its own links initiates recovery by signalling the TM (the opposite node) to coordinate the 2PC across the surviving links, using OAE’s successively reversible protocols to return the state machines on both sides to quiescence—ensuring that no data structures are left stale or inconsistent. Because each node is simultaneously a participant on two links and a transaction manager for the third, the triangle is the minimal topology in which every link has a designated recovery coordinator—and that coordinator can always reach both endpoints, since the two surviving links are, by construction, operational.
Gray’s classical 2PC has a well-known blocking vulnerability: if the coordinator fails after prepare but before commit, participants are stranded. In the triangle topology this vulnerability is eliminated: the coordinator for link Alice–Bob is Charlie, who is not on the failed link. Charlie’s failure would be a different failure—detected and recovered by a different triangle in the larger mesh. The failure modes are fully separated: the node that coordinates recovery for a failed link is, by topological construction, not a victim of that failure.
10.2.1 Scale Independence without Switches
The triangle is only a fragment of a larger structure. Figure 3 shows how triangles compose into an octavalent substrate: a rectangular array of nodes where each node has eight ports connecting to its neighbours in the four cardinal directions (N, S, E, W) and the four diagonal directions (NE, NW, SE, SW). Every adjacent pair of nodes shares a direct link, and every set of three mutual neighbours forms a triangle with a designated TM for each edge.
This geometry introduces scale independence without switches. The planar array extends indefinitely in all four cardinal directions; the diagonal links provide the triangles required for failover at every point. Adding a row or column adds nodes and links but introduces no new switches, no new shared-medium contention points, and no new categories of silent failure. The topology scales while preserving the causal guarantees of every constituent triangle [32].
Figure 3: A $3\times 3$ octavalent substrate. Each XPU connects to its neighbours in all eight directions (N, S, E, W and all four diagonals) via direct AELinks. Every adjacent triple forms a triangle with a designated TM for each edge. The array extends indefinitely without switches, providing scale independence. From [32].
Applied to iCloud: the system has no triangle topology at all. Each device connects to a central cloud server through the public internet—a switched network with unbounded intermediate hops. There is no direct link between devices, no bilateral state, no mechanism for one device to learn about another’s state except through the server’s eventual-consistency model. The architecture is structurally incapable of maintaining the shared state that correct synchronization requires.
But it could be. Apple users routinely have multiple devices in the same household—MacBooks, iMacs, iPhones, iPads—all within direct wireless reach of one another via Wi-Fi, Bluetooth, or even Thunderbolt cable. These devices already form a physical cluster capable of supporting triangle topologies. There is no fundamental reason that household networks could not adopt OAE’s transactional link protocols, enabling small clusters of Apple devices to maintain bilateral state, detect failures locally, and recover via triangle failover—all without routing through a distant cloud server. The opportunity for Apple to transform iCloud from an eventually-consistent cloud relay into a resilient local mesh, augmented by cloud backup, is both technically feasible and architecturally clear.
11  Link Flapping and the Destruction of Topology Knowledge
The iCloud failure pattern is not confined to file synchronization. The same structural error—FITO semantics amplifying transient uncertainty into systemic failure—appears in network fabrics as link flapping.
11.1  The Scaling Reality
If a single link exhibits a mean time to flap (MTTF) of $T$, then a system with $N$ independent links experiences a flap approximately every:
$T_{\text{cluster}}\approx\frac{T}{N}$
At modern AI scale (hundreds of thousands of optical and electrical links), even rare transient disturbances become a continuous background process [3].
11.2  The Three-State Collapse
At the moment of disturbance, the system may be in one of three states:
  1. 1.
    The data was delivered.
  2. 2.
    The data was not delivered.
  3. 3.
    The system cannot currently know which is true.
FITO protocols collapse all three into a single category: error. Timeouts, retries, and resets are triggered not by known failure, but by unresolved uncertainty. This is the same epistemic collapse that iCloud exhibits when it cannot determine which version of a file is authoritative and resorts to numbered suffixes.
11.3  Failure Amplification
The collapse produces a characteristic loop:
  1. 1.
    Transient physical disturbance introduces uncertainty.
  2. 2.
    Timeout fires.
  3. 3.
    Retries increase traffic and noise.
  4. 4.
    Error rates rise further.
  5. 5.
    Links retrain or reset.
  6. 6.
    Global state is disrupted.
Each step is locally rational. The global effect is instability that far exceeds the underlying physical disturbance. This is structurally identical to iCloud’s failure amplification: a single concurrent edit produces duplicate directory trees, orphaned refs, and cascading conflicts that far exceed the original divergence.
11.4  Topology Knowledge Is Information
When a link flaps, the network loses information about its own topology. Routing tables, spanning trees, and forwarding state depend on knowledge of which links are up. A flap forces the system to re-derive this knowledge—but the re-derivation itself generates traffic, which stresses marginal links, which causes more flaps. Information is destroyed faster than it can be reconstructed.
This is the same structure as iCloud’s FSEvents circular dependency with Time Machine: the system that tracks state changes is destabilized by the process of tracking state changes.
12  The Graph-Theoretic Framework
Distributed systems are trees on top of DAGs on top of graphs [4]:
Graphs:
The physical fabric. Nodes are compute units; edges are communication links with asymmetric bandwidth, latency, and failure characteristics.
DAGs:
Causality, scheduling, and locking. Operations must follow directed partial orders to preserve consistency.
Trees:
Namespace hierarchies, leadership structures, commit chains. Authority is delegated through tree structures that impose total order on specific subsets of the DAG.
iCloud’s Category Mistake is the collapse of this layered structure. The sync engine operates on a DAG (the distributed event graph of uploads, downloads, merges, and conflict resolutions). Time Machine and git operate on trees (the filesystem namespace; the commit graph). The filesystem facade pretends to be a tree when the underlying reality is a DAG. The resulting projection:
$F:\text{DAG (distributed sync state)}\rightarrow\text{Tree (local filesystem)}$
is not faithful. Information about concurrency, conflict sets, device-of-origin, and causal lineage is destroyed.
13  The OAE Resolution
Open Atomic Ethernet (OAE) is designed to provide the structural foundation for resolving the class of failures described above. The resolution is not a “fix” for iCloud. It is a demonstration that protocol semantics exist which avoid the Category Mistake, and that these semantics can be implemented.
13.1  The Core Principle
A transfer is not real until it is known to be real.
OAE enforces perfect information feedback at the link boundary. Data is only considered delivered once it has been explicitly reflected back to the sender. Knowledge, not elapsed time, defines correctness [3].
13.2  Bilateral Transactions
OAE links are not passive channels. They are joint stateful systems. Both peers implement identical state machines that evolve synchronously via the exchange of fixed-size, causally significant tokens. There is no concept of master/slave. Either side may assume the role of initiator or responder, depending on who possesses the token [5].
This is designed to eliminate the class of ambiguity that causes iCloud’s failures:
  • No speculative forward progress (no FITO).
  • No timeout-and-retry (uncertainty pauses progress rather than corrupting state).
  • No epistemic collapse (the three states—delivered, not delivered, unknown—remain distinguished).
13.3  Symmetric Reversibility
For every operation on an OAE link, there exists a logically defined inverse that restores the prior state [6]:
  • Partial transactions can be aborted cleanly, returning to equilibrium with no partial state leaked on either side.
  • Errors (bit flips, packet loss) can be rolled back without corrupting state.
  • All token transfers are atomic: they either complete fully or leave the system unchanged.
Applied to file synchronization, this means: a sync operation that encounters uncertainty does not produce a numbered suffix, a duplicate directory, or a silently overwritten file. It reverses to the last known-good state and waits for resolution. Recovery preserves conserved invariants rather than replaying timestamps.
13.4  Conservation of Information
OAE treats every token as a conserved object: inserted, tracked, and retired without ambiguity. No token is ever “lost”—it is either delivered or explicitly rejected. This is the Conservation Principle applied to networking: the total information in the system is accounted for at all times.
Contrast with iCloud: when a file is modified concurrently on two devices, the information about both modifications must be preserved. iCloud’s last-writer-wins destroys the losing modification. OAE’s bilateral protocol ensures that both sides know about the divergence before either side commits, enabling semantic reconciliation rather than timestamp arbitration.
13.5  Surviving Link Flapping
Because OAE links maintain bilateral state, a transient disturbance does not destroy topology knowledge. When a link flaps:
  1. 1.
    Both sides detect the disturbance through the absence of the expected token.
  2. 2.
    The link enters a recovery state (not an error state).
  3. 3.
    The bilateral state machine determines what was in flight at the moment of disturbance.
  4. 4.
    Any in-flight tokens are either confirmed delivered or reversed.
  5. 5.
    Normal operation resumes from a known state, not a guessed state.
Under this model, the failure is contained within the link’s own failure domain. No routing table invalidation is required. No spanning tree recalculation is triggered. No retry storm propagates. These are design properties of the OAE link protocol; formal proofs and quantitative validation are subjects of ongoing work.
14  Implications for a Correct Sync Model
The preceding analysis implies requirements for any system that claims to synchronize files correctly across distributed devices:
  1. 1.
    Correctness over state, not event history. Reconciliation must operate on invariants, not timestamps.
  2. 2.
    Application-level atomicity preserved end-to-end. Multi-file save sequences must not leak intermediate states.
  3. 3.
    Concurrency explicit and semantically handled. Concurrent modifications must be surfaced as a structured divergence, not arbitrated by clocks.
  4. 4.
    Recovery by invariant, not by timestamp. Restoring “the state at time tt” is insufficient; the system must restore conserved invariants.
  5. 5.
    Audit and rollback via causal structure. The event graph must be preserved, not projected onto a linear chain.
  6. 6.
    Bilateral confirmation before commitment. No modification is considered “synced” until both sides confirm receipt and consistency.
  7. 7.
    Conservation of information. Every file version, every modification, every conflict must be accounted for. Silent destruction of data is a protocol violation.
These are not aspirational. They are the properties that OAE is designed to provide at the link layer. Extending them to file synchronization is an engineering problem, not a theoretical impossibility.
15  Summary
iCloud Drive fails—with Time Machine, with git, with Claude Code, with any tool that assumes POSIX filesystem semantics—because it commits the same Category Mistake that pervades distributed computing: projecting a distributed causal graph onto a linear temporal chain and treating the projection as authoritative.
The failures are not random bugs. They are projection artifacts—predicted by Parker et al. in 1983 and confirmed by four decades of operational experience:
  • Time Machine captures a snapshot of a moving target and calls it history.
  • Git’s lock files become propagation vectors for corruption.
  • Automated toolchains read dataless files that stat() says exist but contain no data.
  • Extended attributes are silently stripped, destroying metadata invariants.
  • Link flapping amplifies transient uncertainty into systemic collapse.
  • Years of normal use produce hundreds of gigabytes of unresolvable divergence, requiring custom tooling to repair.
Open Atomic Ethernet is designed to demonstrate that these failures are not inevitable. By replacing FITO semantics with bilateral, reversible, conservation-preserving transactions, OAE aims to align protocol behavior with physical reality: uncertainty pauses progress rather than corrupting state; recovery restores known invariants rather than replaying guessed timelines; information is conserved, not destroyed.
The question is not whether iCloud can be fixed. The question is whether we are willing to recognize that the Category Mistake—treating cloud sync as a filesystem, treating partial order as total order, treating uncertainty as error—is the source of the problem, and that the solution has been available all along in the structure of physics itself.
Acknowledgements
The Mulligan Stew Gang
The ideas in this document were sharpened over 130+ consecutive weeks of Friday morning sessions with the Mulligan Stew Gang, a working group that evolved from ItsAboutTime.club. The members—physicists, mathematicians, computer scientists, network architects, and practicing engineers—wish to remain anonymous.
Appendix A Observed Failure Log
This appendix documents specific iCloud Drive failures observed during the preparation of this document and related work. Each incident is recorded with dates, error output, and the resolution applied. The intent is to provide a growing evidence base that can be extended as new failures are encountered.
A.1  Incident 1: LaTeX + Git Collaboration Corruption
Date: 02026-FEB-22. Context: Collaborative LaTeX manuscript tracked in Git, stored in iCloud Drive.
During multi-author editing, iCloud’s sync daemon propagated intermediate file states to other devices. Figure paths became invalid as iCloud renamed files during sync. Git operations failed with spurious merge conflicts that had no basis in actual concurrent edits [26].
Resolution: The entire repository was moved out of iCloud Drive to a local directory. Collaboration continued via Git remote, with no further corruption.
A.2  Incident 2: Silent Filename Swap
Date: 02023-MAY (approx). Context: Creating a new LaTeX document by copying an existing file in an iCloud-synced directory.
iCloud silently swapped filenames during a copy operation. A LaTeX source comment records the event directly: “I edited Subtime-copy.tex to create this—iCloud should leave it alone and not swap the filenames back” [27]. The sync daemon treated a deliberate file creation as a conflict requiring resolution, and “resolved” it by undoing the user’s work.
Resolution: Manual correction. The comment was left in the source as a warning.
A.3  Incident 3: Repository Seizure During git init
Date: 02026-FEB-22. Context: Attempting to create a Git repository in the iCloud-synced directory containing this paper’s source files (four files, under 200 KB total, no concurrent access from any other device).
When git init was run, the sync daemon immediately seized the newly created .git/ directory. The first git commit failed:
fatal: Unable to create ’.git/index.lock’: File exists.
Attempting to remove the stale lock file:
rm: cannot remove ’.git/index.lock’: Operation not permitted
Attempting to remove the entire .git/ directory:
rm: cannot remove ’.git/index.lock’: Operation not permitted
rm: cannot remove ’.git/config’: Operation not permitted
rm: cannot remove ’.git/HEAD’: Operation not permitted
rm: cannot remove ’.git/hooks/commit-msg.sample’: Operation not permitted
rm: cannot remove ’.git/hooks/pre-commit.sample’: Operation not permitted
rm: cannot remove ’.git/refs/heads’: Operation not permitted
rm: cannot remove ’.git/refs/tags’: Operation not permitted
    [... 24 files total, every one "Operation not permitted"]
The sync daemon had acquired exclusive access to every file in the .git/ directory and would not release any of them. The repository was permanently inoperable within seconds of its creation.
Resolution: The repository was initialized outside iCloud Drive, source files were copied in, and the resulting .bundle file was placed in the iCloud-synced directory for archival—exactly the workaround documented in Section 5. The paper that documents the incompatibility could only be version-controlled by routing around it. The trapped .git/ directory had to be deleted manually from Finder on the host Mac.
A.4  Incident 4: Keynote Package Image Degradation
Date: 02026-FEB-22. Context: Opening a Keynote presentation stored in iCloud Drive.
Upon opening DAE-Investor-Primary copy.key—note the “copy” suffix indicating iCloud had previously created a conflict duplicate—Keynote displayed a warning dialog:
“This presentation contains images that may be damaged and appear in low resolution. Would you like to open the file anyway?”
Keynote files are macOS package bundles: directories containing XML metadata, theme data, and embedded media files in an internal structure with cross-references between components. iCloud syncs the component files independently rather than treating the package as an atomic unit. When timing or fidelity of the sync breaks the internal consistency of the package, embedded resources—in this case, presentation images—are degraded or lost.
This failure demonstrates a distinct incompatibility axis from the git and filesystem failures: application bundle formats that depend on internal structural consistency are vulnerable to iCloud’s file-by-file sync model. The same class of failure can affect any macOS package format, including .pages, .numbers, Xcode projects (.xcodeproj), and Core Data stores.
Resolution: No automated resolution available. The damaged images would need to be re-inserted manually from original sources, if they still exist elsewhere.
A.5  Incident 5: Keynote File Stuck on Open
Date: 02026-FEB-22. Context: Attempting to open a Keynote presentation stored in iCloud Drive in order to export it as PDF.
While attempting to open Spec 0.7A release.key from the PRESENTATIONS/OAE-Presentations/ directory, Keynote displayed a progress bar that froze indefinitely. The file could not be opened or exported. This occurred immediately after Incident 4 (image degradation in a different Keynote file), suggesting that iCloud’s sync state for Keynote package bundles was in a degraded condition.
This failure demonstrates the dataless file problem at the application level: the file metadata was present (Finder listed it, Keynote attempted to open it), but the actual data extents—the internal package contents—were either absent, partially hydrated, or corrupted during on-demand download. Keynote’s progress bar represented the application waiting for data that iCloud could not deliver.
Resolution: None during the session. The file remained inaccessible.